OpenStudy (anonymous):

Use the quadratic equation to solve 3x2 + x – 2 = 0 for x exactly.

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] does that helps

OpenStudy (anonymous):

whats the quadratic equation? im kind of confused.

OpenStudy (anonymous):

ax^2+bx+c=0 x= the equation i just sent

OpenStudy (anonymous):

want to try